OpenStudy (anonymous):

A 776 N student stands in the middle of a frozen pond having a radius of 4.2 m. He is unable to get to the other side because of a lack of friction between his shoes and the ice. To overcome this difficulty, he throws his 2.7 kg physics textbook horizontally toward the north shore at a speed of 5.3 m/s. How long does it take him to reach the south shore?

OpenStudy (cwrw238):

i think you can do this using conservation of momentum the students momentum will equal that of of the book so we have 776 * v = 2.7 * 9.8 * 5.3 v = (2.7*9.8 *5.5) / 776 m / s then you can find the time taken